# Experience of Training a 1.7B-Parameter LLaMa Model From Scratch

### v2

@@ III-B Qualitative Evaluation of Generated Text
By checkpoint 35000, the text becomes significantly more coherent and contextually relevant, though it introduces minor inaccuracies, such as the association with Henry VIII. At checkpoint 43250, the output demonstrates marked improvement, correctly identifying William I as ”William The Conqueror” and providing accurate dates and titles.
These observations align with the scores shown in Table II: less sufficiently trained language models are more prone to “hallucination”—the production of factually incorrect or contextually inconsistent details—while more extensively trained models show a marked reduction in such errors. As the model’s training proceeds, it appears to acquire a stronger grounding in historical context, produce more thematically coherent narratives, and adhere more closely to factual accuracy. In other words, increased training not only enhances fluency and coherence but also mitigates the tendency to fabricate details, resulting in outputs that are both stylistically refined and factually sound.
TABLE IV: Validation Loss and Hellaswag Scores Across Training Steps
@@
In addition to subjective assessment, we further analyzed the relationship between the average qualitative score and three key variables: training steps, validation loss, and Hellaswag [13] scores (see Table IV for the progression of these metrics). Correlation analysis yielded the following results:
III-C Comparison and Discussion
The performance of DMaS-LLaMa-Lite is compared against TinyLLaMa’s 2T checkpoint [14], a series of lightweight LLaMa-based models trained with significantly larger token counts. TinyLLaMa’s 2T checkpoint, a 1.1B-parameter model trained on 2 trillion tokens using the SlimPajama dataset [15], represents an established baseline for small-scale LLMs.
The benchmarks used for evaluation are as follows:
  1. 1.
    ARC Challenge: A subset of the AI2 Reasoning Challenge [16], which focuses on grade-school science questions that require reasoning and commonsense knowledge.
  2. 2.
  3. 3.
    BoolQ: A yes/no question-answering benchmark [17] that requires reading comprehension and contextual reasoning.
  4. 4.
    HellaSwag: A commonsense reasoning benchmark [13] where the goal is to select the most plausible continuation of a given sentence.
  5. 5.
    OpenBookQA: A multiple-choice question benchmark [18] that requires reasoning over a small “open book” of facts.
  6. 6.
    PIQA: The Physical Interaction QA benchmark [19], which tests commonsense physical reasoning, such as how objects are used.
  7. 7.
    Winogrande: A large-scale Winograd Schema Challenge [20] dataset, focusing on resolving pronoun ambiguity using commonsense reasoning.
TABLE V: Comparison of Responses Before and After Post-Training Instruction Tuning
@@ III-C Comparison and Discussion
A notable exception to the general trend is the BoolQ benchmark [17]. Our model’s performance not only lags behind TinyLLaMa at 2T tokens but also diminishes slightly as training progresses. Intriguingly, this degradation in BoolQ performance parallels a similar pattern observed in TinyLLaMa: while TinyLLaMa initially achieves a BoolQ accuracy of around 63.21% at 2T tokens, its accuracy declines to 57.83% by the time it reaches 3T tokens of training. This phenomenon suggests a shared difficulty that may not stem from data quality or scale alone but from the intrinsic complexity of BoolQ-style yes/no question answering.
Passage: ”Windows Movie Maker (formerly known as Windows Live Movie Maker in Windows 7) is a discontinued video editing software by Microsoft. It is a part of Windows Essentials software suite and offers the ability to create and edit videos as well as to publish them on OneDrive, Facebook, Vimeo, YouTube, and Flickr.”
Question: ”Is Windows Movie Maker part of Windows Essentials?”
Answer: ”Yes”
This subtlety underscores that BoolQ tasks require interpretable reasoning. Yet, the benchmark format itself does not give the model an opportunity to produce an explicit chain-of-thought or reasoning trace to use test time computation [21], but rather the task is framed to predict ”yes” or ”no” based on the passage content. Without such a mechanism, the model must supply an immediate yes/no answer, leaving no room to display intermediate reasoning steps that might help ensure correctness. Thus, the challenge is twofold: models need robust reasoning capabilities, and the benchmark structure (which does not request or evaluate reasoning traces) limits the model’s ability to leverage explicit reasoning strategies that could improve performance.
IV Posttraining
Following the initial pretraining phase, we conducted a post-training step to refine the model’s capacity for following user instructions and producing contextually appropriate, helpful responses. This stage leverages techniques akin to instruction tuning, where a base model—already competent in general text generation—is further specialized to respond more effectively to prompts in a conversational format.
IV-A Instruction-Following Data and Template
For the post-training stage, we used the yahma/alpaca-cleaned  [^3] https://huggingface.co/datasets/yahma/alpaca-cleaned dataset, a widely utilized instruction-tuning corpus derived from the Stanford Alpaca dataset [22]. To present this data to the model in a consistent, conversational manner, we applied a Vicuna 1.1-style chat template [23], which structures the interaction as a user-assistant conversation:
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: input
ASSISTANT: output<|endoftext|>
Here, input is replaced with the user’s query or instruction, and output with the model’s expected answer. The ¡—endoftext—¿ token serves as a signal for the model to stop generating text at the appropriate point.
IV-B Model Initialization and Parameter-Efficient Fine-Tuning
We began the post-training step from the 43,500-step checkpoint of DMaS-LLaMa-Lite. Rather than fine-tuning all parameters, we employed a parameter-efficient technique based on Low-Rank Adaptation (LoRA) [24]. We injected LoRA adapters into attention-related parameters (e.g., q_proj, k_proj, v_proj, o_proj, and feed-forward projections) using a rank of $r=128$, a LoRA $\alpha=16$, and no dropout. This approach reduces the memory footprint and computational overhead while enabling efficient adaptation of the model to the instruction-following domain.
IV-C Training Dynamics
@@
IV-D Qualitative Improvements in Responses
Table V compares the conversational behaviors of the models before and after post-training using two example prompts. Interestingly, the model demonstrates an ability to answer questions even prior to post-training. This phenomenon can be attributed to two factors. First, the Vicuna 1.1 prompting format aligns closely with natural conversational scenarios and clearly indicates that the assistant is expected to provide answers. Consequently, the model can generate responses based on its pretraining alone, without requiring additional fine-tuning. Second, it is possible that the pretraining dataset includes instruction-tuning samples formatted similarly to Vicuna 1.1 prompts, as this format predates the pretraining corpus’s knowledge cutoff.
Another noteworthy observation is that instruction tuning during post-training significantly enhances the detail and contextual depth of the model’s responses. Compared to the pretraining-only model, the post-trained version consistently generates more elaborate and contextually rich answers, reflecting improved understanding and engagement with the prompts.
IV-E Sensitivity to Prompt Variations
@@
One plausible explanation involves the underlying tokenizer and how it handles spaces. In GPT-2-style tokenizers, spaces are generally prefixed to words rather than suffixed. This leads to subtle differences in the tokenization of seemingly similar inputs. For example, for an answer in the training set: ”ASSISTANT: Berlin was founded in the 13th century.”, it is tokenized as: [”ASS”, ”IST”, ”ANT”, ”:”, ”ĠBerlin”, ”Ġwas”, ”Ġfounded”, ”Ġin”, ”Ġthe”, ”Ġ13”, ”th”, ”Ġcentury”, ”.”]. Here, “Ġ” represents a space token. If the prompt ends with ”ASSISTANT: ”, however, the tokenization changes to: [”ASS”, ”IST”, ”ANT”, ”:”, ”Ġ”]. The inclusion of the extra ”Ġ” token alters the model’s understanding of the input and significantly shifts the probability distribution for the next token. This is due to the statistical co-occurrence patterns learned during pretraining, which differ depending on the presence of the additional space token.
@@
V Conclusion
In this paper, we documented the pretraining experience of DMaS-LLaMa-Lite, a fully open-source, 1.7-billion-parameter LLaMa-based language model trained on approximately 20 billion tokens of a carefully curated dataset. By chronicling the model’s training trajectory through validation loss, downstream benchmarks, qualitative evaluations, and practical challenges, we demonstrated how the training process influences model outputs and performance.
@@ V Conclusion
Our findings, along with the provided training logs, checkpoints, and sample outputs, aim to guide future researchers and practitioners in refining both pretraining and post-training strategies. By documenting the full trajectory from pretraining dynamics to post-training refinement, we emphasize a process-focused perspective where transparency and detailed insights foster reproducibility and more effective model development. This approach lays a foundation for future studies to explore additional best practices, methodological refinements, and novel techniques in both pretraining and instruction-tuning phases of large language models.

### v3

@@ 3.3 Comparison and Discussion
A notable exception to the general trend is the BoolQ benchmark (Clark et al., 2019). Our model’s performance not only lags behind TinyLLaMa at 2T tokens but also diminishes slightly as training progresses. Intriguingly, this degradation in BoolQ performance parallels a similar pattern observed in TinyLLaMa: while TinyLLaMa initially achieves a BoolQ accuracy of around 63.21% at 2T tokens, its accuracy declines to 57.83% by the time it reaches 3T tokens of training. This phenomenon suggests a shared difficulty that may not stem from data quality or scale alone but from the intrinsic complexity of BoolQ-style yes/no question answering.
Passage: "Windows Movie Maker (formerly known as Windows Live Movie Maker in Windows 7) is a discontinued video editing software by Microsoft. It is a part of Windows Essentials software suite and offers the ability to create and edit videos as well as to publish them on OneDrive, Facebook, Vimeo, YouTube, and Flickr."
Question: "Is Windows Movie Maker part of Windows Essentials?"
Answer: "Yes"
This subtlety underscores that BoolQ tasks require interpretable reasoning. Yet, the benchmark format itself does not give the model an opportunity to produce an explicit chain-of-thought or reasoning trace to use test time computation (Snell et al., 2024), but rather the task is framed to predict "yes" or "no" based on the passage content. Without such a mechanism, the model must supply an immediate yes/no answer, leaving no room to display intermediate reasoning steps that might help ensure correctness. Thus, the challenge is twofold: models need robust reasoning capabilities, and the benchmark structure (which does not request or evaluate reasoning traces) limits the model’s ability to leverage explicit reasoning strategies that could improve performance.
Table 4: Validation Loss and Hellaswag Scores Across Training Steps
@@
4 Post-training
Following the initial pre-training phase, we conducted a post-training step to refine the model’s capacity for following user instructions and producing contextually appropriate, helpful responses. This stage leverages techniques akin to instruction tuning, where a base model—already competent in general text generation—is further specialized to respond more effectively to prompts in a conversational format.
4.1 Instruction-Following Data and Template
For the post-training stage, we used the yahma/alpaca-cleaned  [^5] https://huggingface.co/datasets/yahma/alpaca-cleaned dataset, a widely utilized instruction-tuning corpus derived from the Stanford Alpaca dataset (Taori et al., 2023). To present this data to the model in a consistent, conversational manner, we applied a Vicuna 1.1-style chat template (Zheng et al., 2023), which structures the interaction as a user-assistant conversation:
A chat between a curious user and an artificial intelligence assistant. The assistant
gives helpful, detailed, and polite answers to the user’s questions.
USER: input
ASSISTANT: output<|endoftext|>
Here, input is replaced with the user’s query or instruction, and output with the model’s expected answer. The <|endoftext|> token serves as a signal for the model to stop generating text at the appropriate point.
4.2 Model Initialization and Parameter-Efficient Fine-Tuning
We began the post-training step from the 43,500-step checkpoint of DMaS-LLaMa-Lite. Rather than fine-tuning all parameters, we employed a parameter-efficient technique based on Low-Rank Adaptation (LoRA) (Hu et al., 2021). We injected LoRA adapters into attention-related parameters (e.g., q_proj, k_proj, v_proj, o_proj, and feed-forward projections) using a rank of $r=128$, a LoRA $\alpha=16$, and no dropout. This approach reduces the memory footprint and computational overhead while enabling efficient adaptation of the model to the instruction-following domain.
4.3 Training Dynamics
Table 5: Comparison of Responses Before and After Post-Training Instruction Tuning
@@ 4.3 Training Dynamics
4.4 Qualitative Improvements in Responses
Table 5 compares the conversational behaviors of the models before and after post-training using two example prompts. Interestingly, the model demonstrates an ability to answer questions even prior to post-training. This phenomenon can be attributed to two factors. First, the Vicuna 1.1 prompting format aligns closely with natural conversational scenarios and clearly indicates that the assistant is expected to provide answers. Consequently, the model can generate responses based on its pre-training alone, without requiring additional fine-tuning. Second, it is possible that the pre-training dataset includes instruction-tuning samples formatted similarly to Vicuna 1.1 prompts, as this format predates the pre-training corpus’s knowledge cutoff.
Another noteworthy observation is that instruction tuning during post-training significantly enhances the detail and contextual depth of the model’s responses. Compared to the pre-training-only model, the post-trained version consistently generates more elaborate and contextually rich answers, reflecting improved understanding and engagement with the prompts.
4.5 Sensitivity to Prompt Variations
@@
One plausible explanation involves the underlying tokenizer and how it handles spaces. In GPT-2-style tokenizers, spaces are generally prefixed to words rather than suffixed. This leads to subtle differences in the tokenization of seemingly similar inputs. For example, for an answer in the training set: "ASSISTANT: Berlin was founded in the 13th century.", it is tokenized as: ["ASS", "IST", "ANT", ":", "ĠBerlin", "Ġwas", "Ġfounded", "Ġin", "Ġthe", "Ġ13", "th", "Ġcentury", "."]. Here, “Ġ” represents a space token. If the prompt ends with "ASSISTANT: ", however, the tokenization changes to: ["ASS", "IST", "ANT", ":", "Ġ"]. The inclusion of the extra "Ġ" token alters the model’s understanding of the input and significantly shifts the probability distribution for the next token. This is due to the statistical co-occurrence patterns learned during pre-training, which differ depending on the presence of the additional space token.
@@
5 Conclusion
In this paper, we documented the pre-training experience of DMaS-LLaMa-Lite, a fully open-source, 1.7-billion-parameter LLaMa-based language model trained on approximately 20 billion tokens of a carefully curated dataset. By chronicling the model’s training trajectory through validation loss, downstream benchmarks, qualitative evaluations, and practical challenges, we demonstrated how the training process influences model outputs and performance.
@@ 5 Conclusion
Our findings, along with the provided training logs, checkpoints, and sample outputs, aim to guide future researchers and practitioners in refining both pre-training and post-training strategies. By documenting the full trajectory from pre-training dynamics to post-training refinement, we emphasize a process-focused perspective where transparency and detailed insights foster reproducibility and more effective model development. This approach lays a foundation for future studies to explore additional best practices, methodological refinements, and novel techniques in both pre-training and instruction-tuning phases of large language models.
Acknowledgements
The research is supported in part by Discovery Grants (RGPIN-2024-04087) from the Natural Sciences and Engineering Research Council of Canada (NSERC), Canada Research Chairs Program (CRC-2019-00041), and R&D Spearhead Grants (2024-1301) from the National Cybersecurity Consortium (NCC).